40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Dec 18 11:37:12 2013, Jenne, Update, LSC, POP22 and POP110 demod phases 
    Reply  Wed Dec 18 18:45:39 2013, Jenne, Update, LSC, REFL 165 demod phases 
       Reply  Mon Jan 6 22:11:46 2014, Jenne, Update, LSC, PRCL sideband locking still not so happy 
Message ID: 9491     Entry time: Wed Dec 18 18:45:39 2013     In reply to: 9487     Reply to this: 9523
Author: Jenne 
Type: Update 
Category: LSC 
Subject: REFL 165 demod phases 

I checked out the REFL165 demod phase, and it looks like it was okay.  it was -20.9 degrees.  I turned on my sensing matrix oscillators, and maximized the PRCL signal in the REFL165_I_ERR channel, and got a pretty good maximization at +155 degrees.  I used this to lock the PRMI on sidebands, with MICH gain of +0.3, and PRCL gain of +0.1 . 

Since this is working, I'm leaving the REFL 165 phase here, at +155 degrees, although this is almost exactly 180 degrees from what Den left it at, so I'm not sure why I was not able to lock with a demod phase of -20.9.  (I tried all 4 permutations of signs, with gain values of the same magnitude (0.3 for MICH, 0.1 for PRCL), and wasn't able to lock.  I'll try to figure this out tomorrow, but it was time for the meeting, then the IFO has been busy doing more important things the rest of the afternoon.

Plan for checking:  Lock with demod phase of 155, measure TF to one of the other REFL diodes (11, 33 or 55), lock on that other REFL diode.  Then, change the REFL 165 demod phase back to -20.9, and measure the transfer function again.  Hopefully the answer is just that I was doing something dumb, and it works easily.  This test/measurement should only take a few minutes, but it'll make me happier knowing that things still work as they should.

ELOG V3.1.3-